# Positive and Unlabeled Data: Model, Estimation, Inference, and Classification

### v2

@@ Abstract
This study introduces a new approach to addressing positive and unlabeled (PU) data through the double exponential tilting model (DETM). Traditional methods often fall short because they only apply to selected completely at random (SCAR) PU data, where the labeled positive and unlabeled positive data are assumed to be from the same distribution. In contrast, our DETM’s dual structure effectively accommodates the more complex and underexplored selected at random PU data, where the labeled and unlabeled positive data can be from different distributions. We rigorously establish the theoretical foundations of DETM, including identifiability, parameter estimation, and asymptotic properties. Additionally, we move forward to statistical inference by developing a goodness-of-fit test for the SCAR condition and constructing confidence intervals for the proportion of positive instances in the target domain. We leverage an approximated Bayes classifier for classification tasks, demonstrating DETM’s robust performance in prediction. Through theoretical insights and practical applications, this study highlights DETM as a comprehensive framework for addressing the challenges of PU data.
@@ 1.1 Positive and Unlabeled Data
Considering a binary response variable $Y\in\{0,1\}$ and features ${\bf X}\in\mathbb{R}^{p}$, positive and unlabeled (PU) data refer to datasets that consist of the following two samples: The first sample is labeled, but all observations have positive responses $y=1$; the second sample contains both positive ($y=1$) and negative ($y=0$) instances. However, the second sample is unlabeled, meaning we can only access the features ${\bf x}$ but have no access to their labels $y$. PU data are prevalent in many areas, including biomedical study, ecology, and machine learning.
@@
A case-control study is a research design widely used in biomedical research to investigate the causes of diseases. It compares two groups: a group of patients with the disease (the case group) and a group of healthy people (the control group). Data contamination is a common issue in case-control studies. For example, the underdiagnosis phenomenon refers to a situation where the false positive rate is very low, and the case group can generally be seen as pure and only contains patients. However, some medical conditions may go unrecognized in many patients, and they are misclassified as healthy ones. Then, many people in the control group are, in fact, patients. For example, Begg, (1994) and Godley and Schell, (1999) highlighted an issue in a prostate cancer study, indicating that 20%-40% of the control group consists of latent prostate cancer cases, rendering the control group contaminated. In the underdiagnosis example, the case group can be seen as a positive-only sample, but the contaminated control group is an unlabeled sample containing both positive and negative cases because we do not know who are real healthy people and patients.
@@
Presence-only data can come from various fields, including ecology, epidemiology, and marketing. For example, ecological studies often rely on presence-only data to model a species’ habitat (Ward et al., 2009). The dataset comprises two samples: one contains confirmed presence locations of a species obtained through field surveys or other methods; the other contains locations sampled from the full study area. However, the presence status of the species in the latter sample is unknown. Thus, the first sample is positive only while the second sample contains both presence and non-presence locations but is unlabeled, meaning that we do not know which location has the confirmed presence of the species.
@@
PU data are commonly seen in many machine learning tasks. For instance, on an online video streaming platform, users click on videos that they find interesting. However, for videos that are not clicked, it is uncertain whether users are interested in them or not, as users may ignore those videos while browsing. In such a scenario, those clicked videos are positive-only data, while unclicked videos are considered unlabeled data for a recommendation system. In computer vision, tasks like facial recognition are often based on PU data. Take this specific example: We aim to identify a user from various images using some known reference pictures of that person. Here, the labeled sample comprises all clearly identified face images of that user. The unlabeled sample we wish to identify consists of photos of the user and others. We refer interested readers to read Jaskie and Spanias, (2019) for more examples of PU data in machine learning.
@@ 1.2 Notation and Motivations
Even though much research has been done on PU data, most of it suffers from two major limitations. The first limitation is that many existing works require the mixture proportion $\pi$ to be known. However, such an assumption is usually unrealistic as $\pi$ is rarely known. Even with prior knowledge of mixture proportion, one must use an estimate to replace the underlying true value. However, it is generally unclear how the difference between the estimated and true values will affect the analysis results.
The second major limitation is that most existing research works are constrained to the selected completely at random (SCAR) PU data (Bekker and Davis, 2020), which imposes the assumption that $p_{s}({\bf x}|y=1)=p_{t}({\bf x}|y=1)$. The SCAR assumption has been widely used for simplicity: one of the mixture components, $p_{t}({\bf x}|y=1)$, is readily identifiable because the positive-only data $\left\{({\bf x}_{i},y_{i}=1),i=1,\dots,n\right\}$ are from the same distribution as $p_{t}({\bf x}|y=1)$. Despite its popularity, the SCAR assumption does not always reflect the true underlying data-generating mechanism. For example, in the contaminated case-control study example, it is reasonable to believe that the latent prostate cancer cases were misdiagnosed as healthy ones not only because of bad luck but also because their disease symptoms were not as severe as patients in the case group (i.e., $p_{s}({\bf x}|y=1)\neq p_{t}({\bf x}|y=1)$).
As opposed to the SCAR assumption, the selected at random (SAR) PU setting (Bekker and Davis, 2020) allows $p_{s}({\bf x}|y=1)$ to be different from $p_{t}({\bf x}|y=1)$, thus it can better reflect the true data-generating mechanism. Figure 1 shows the difference between SCAR and SAR PU data. However, despite its flexibility, the SAR PU setting presents numerous challenges, the foremost being the identifiability of the distributions of unlabeled positive and unlabeled negative data (i.e., $p_{t}({\bf x}|y=1)$ and $p_{t}({\bf x}|y=0)$).
The aim of this paper is to relax the SCAR assumption and model PU data under the SAR setting. In addition, we aim to propose a method that does not rely on prior knowledge of the mixture proportion $\pi$.
Figure 1: We use $R=0$ (red) and $R=1$ (blue), respectively, to denote data from the target and source. In the SCAR setting, the labeled positive (blue triangles) and unlabeled positive (red triangles) are from the same distribution. However, they may be different in the SAR setting.
@@ 1.3 Related Work and Challenges
PU data have been studied under different names in the statistical literature, including contaminated case-control data (e.g., Qin and Lawless, 1994; Lancaster and Imbens, 1996; Qin, 1999; Qin and Liang, 2011; Duan et al., 2020) and presence-only data (e.g., Ward et al., 2009; Song and Raskutti, 2019). However, the aforementioned works are constrained under the SCAR assumption and often rely on knowing the true value of the mixture proportion. One exception is Wang et al., (2021), where they consider the SAR setting. However, they further require a smaller labeled verification sample from the target distribution. Thus, data from their setting cannot be seen as PU data.
In machine learning literature, most works on PU data focus on the SCAR assumption (e.g., Elkan and Noto, 2008; Ramaswamy et al., 2016; Garg et al., 2021). However, a few works try to address the SAR setting. Kato et al., (2019) impose an “invariance of order” assumption, which states that, for observations in the labeled source data, the higher the probability $\Pr_{s}(y=1|{\bf x})$ is, the closer it is to the unlabeled positive distributing $p_{t}({\bf x}|y=1)$. In addition, their method requires the mixture proportion $\pi=\Pr_{t}(y=1)$ to be known. Bekker et al., (2019) propose a method similar to inverse probability weighting, which requires estimating the propensity score function. To ensure identifiability, they assume that the propensity score function needs to depend on fewer features than the predictive model (i.e., the shadow variable assumption in the missing data literature). Coudray et al., (2023) further establish risk bounds based on Bekker et al., (2019). Nevertheless, the shadow variables are generally unknown, let alone their existence, thus reducing the method’s applicability. Gong et al., (2021) follow a similar framework as in Bekker et al., (2019) but do not impose the shadow variable assumption. However, they require a highly restrictive assumption that the propensity score function has a value of less than 0.5 for all observations from the unlabeled data under the logistic regression model. Furmańczyk et al., (2022) use a similar idea as Bekker et al., (2019) and Gong et al., (2021), in which they prove the model identifiability under logistic regression. However, their prediction approach requires that the positive and negative instances in the unlabeled data are separable. In addition, existing works on the SAR PU setting focus solely on classification (i.e., prediction) and do not address statistical inference. Problems like the the goodness-of-fit of the SCAR condition, and inference of the mixture proportion are remained unexplored.
@@ 1.4 Our Contributions and Overview
In this paper, we propose a new model, called the double exponential tilting model (DETM), for the SAR PU data, and prove that it is identifiable. Based on the proposed model, we develop an Expectation-Maximization (EM) algorithm for estimating parameters using the empirical likelihood (EL) method. It is worth noting that our method works even if the mixture proportion is unknown. Lastly, we apply the proposed model to address inference and prediction tasks, including model selection, confidence intervals, hypothesis testing, and classification. The highlights of our contributions are summarized in the following.
  • Our proposed model, based on modeling the density ratios, is identifiable up to label switching. With some further mild assumptions, the model becomes fully identifiable.
@@
  • We investigate the asymptotic behavior of our proposed estimator and propose hypothesis testing procedures for the mixture proportion and the goodness-of-fit of the SCAR condition.
@@
The rest of this paper is organized as follows. Section 2 introduces the proposed DETM. Section 3 discusses how to estimate parameters under the proposed model. Section 4 establishes theoretical results for statistical inference and classification. Section 5 examines the empirical performance of the proposed methods using simulation studies and real data application. We close the paper with concluding remarks given in Section 6. All the technical details and proofs are given in the supplementary materials.
@@ 2.1 Density Ratio is the Key
A commonly used strategy in distributional shift problems (or, broadly speaking, transfer learning) for connecting the source and target domains is to utilize the density ratio. For example, adjusting the covariate shift often relies on estimating the density ratio of the features ${\bf x}$ in the target and source domains (e.g., Shimodaira, 2000; Sugiyama et al., 2007), and adjusting the label shift often requires estimating the density ratio of the response in the target and source domains (e.g., Lipton et al., 2018; Tian et al., 2023; Kim et al., 2024).
The aforementioned works on transfer learning suggest that modeling the density ratios is a promising direction for solving the SAR PU problem, which also involves a source and a target domain and belongs to the broad concept of transfer learning. In fact, examining the Bayes classifier for the target domain also reveals the key role of the density ratios. The Bayes classifier utilizes the posterior probability of the label and is optimal because it minimizes the prediction risk. The Bayes classifier on the target domain is defined as
@@
where $\mathbbm{1}(\cdot)$ is the indicator function. We can rewrite ${\Pr}_{t}(Y=1|{\bf X})$ using the Bayes rule as
@@
We can further write ${\Pr}_{t}(Y=1|{\bf X})$ by dividing both the numerator and denominator by $p_{s}({\bf x}|y=1)$:
@@
Equation (2) implies that the Bayes classifier hinges on the mixture proportion $\pi$ and two density ratios $\Lambda_{1}({\bf x})$ and $\Lambda_{2}({\bf x})$. This finding also points out that it is promising to model the SAR PU data using the density ratios.
@@ 2.2 Identifiability Problem
We have underscored the importance of $\Lambda_{1}({\bf x})$ and $\Lambda_{2}({\bf x})$ in the previous section. However, we will show that if $\Lambda_{1}({\bf x})$ and $\Lambda_{2}({\bf x})$ are unconstrained, $\pi$, $\Lambda_{1}({\bf x})$, and $\Lambda_{2}({\bf x})$ are non-identifiable.
@@ 2.3 Double Exponential Tilting Model
Based on previous discussions, we use the ETM to connect the source and target domains. Rather than using an exponential tilting model for $p_{t}({\bf x},y)/p_{s}({\bf x},y)$ (e.g., Maity et al., 2023), we acknowledge that the target distribution $p_{t}({\bf x})$ is a mixture of $p_{t}({\bf x}|y=1)$ and $p_{t}({\bf x}|y=0)$. Thus, we connect the source sample distribution $p_{s}({\bf x}|y=1)$ with the two mixture components $p_{t}({\bf x}|y=1)$ and $p_{t}({\bf x}|y=0)$ using two ETMs separately. We call our proposed model the DETM, where the details are given below.
@@
The remaining question is to address the identifiability of the DETM. As we have shown in (3), the key to model identifiability lies in $\pi\Lambda_{1}({\bf x})+(1-\pi)\Lambda_{2}({\bf x})$. In the next theorem, we prove that $\pi\Lambda_{1}({\bf x})+(1-\pi)\Lambda_{2}({\bf x})$ is “almost” identifiable and can become fully identifiable with some mild assumptions. The proof is given in the supplementary materials.
@@
Under the DETM in (4), we model the density ratios as $\Lambda_{1}({\bf x})=\exp\left(\alpha_{1}+{\bf x}^{\rm T}{\boldsymbol{\beta}}\right)$ and $\Lambda_{2}({\bf x})=\exp(\alpha_{2}+{\bf x}^{\rm T}{\boldsymbol{\beta}})$, and we let $\pi={\Pr}_{t}(Y=1)$. If $(\pi,\alpha_{1},{\boldsymbol{\beta}}_{1},\alpha_{2},{\boldsymbol{\beta}}_{2})$ and $(\pi^{\ast},\alpha^{\ast}_{1},{\boldsymbol{\beta}}^{\ast}_{1},\alpha^{\ast}_{2},{\boldsymbol{\beta}}^{\ast}_{2})$ satisfy
@@
for any ${\bf x}\in\mathbb{R}^{p}$, then the two sets of parameters are either the same
$(\pi,\alpha_{1},{\boldsymbol{\beta}}_{1},\alpha_{2},{\boldsymbol{\beta}}_{2})=(\pi^{\ast},\alpha^{\ast}_{1},{\boldsymbol{\beta}}^{\ast}_{1},\alpha^{\ast}_{2},{\boldsymbol{\beta}}^{\ast}_{2})$
or subject to label switching
$(\pi,\alpha_{1},{\boldsymbol{\beta}}_{1},\alpha_{2},{\boldsymbol{\beta}}_{2})=(1-\pi^{\ast},\alpha^{\ast}_{2},{\boldsymbol{\beta}}^{\ast}_{2},\alpha^{\ast}_{1},{\boldsymbol{\beta}}^{\ast}_{1}).$
@@
Theorem 1 reveals that the DETM is identifiable up to exchanging the label $Y$. The label switching issue is inherent in the model specification in (4) due to the symmetry of $\Lambda_{1}({\bf x})$ and $\Lambda_{2}({\bf x})$. We can achieve full model identifiability by introducing some reasonable assumptions to solve the label-switching problem. For instance, if positive cases are less frequent in the unlabeled dataset (e.g., misdiagnosis is rare), we can impose the condition $\pi<0.5$, which resolves the label-switching problem. Alternatively, if we believe the labeled positive source distribution $p_{s}({\bf x}|y=1)$ is more similar to the positive target distribution $p_{t}({\bf x}|y=1)$ than the negative target distribution $p_{t}({\bf x}|y=0)$, we can impose an assumption based on some distance measure, such as the Kullback-Leibler (KL) divergence:
$\mathrm{KL}\left\{p_{s}({\bf x}|y=1)\|p_{t}({\bf x}|y=1)\right\}<\mathrm{KL}\left\{p_{s}({\bf x}|y=1)\|p_{t}({\bf x}|y=0)\right\}.$
In the rest of this paper, we assume the DETM is fully identifiable.
@@
The proposed DETM can be applied to both SCAR and SAR PU data because SCAR data can be viewed as a special case of SAR data with additional constraints. Under the SCAR assumption, the distribution of features of positive cases remains the same in the source and target domains. That is
@@
for any ${\bf x}$. We refer to the constrained model for SCAR PU data as the single exponential tilting model (SETM) because we only have the following model assumption:
@@
The SETM is a reduced model of DETM because we consider a reduced parameter space where $\alpha_{1}=0$ and ${\boldsymbol{\beta}}_{1}={\bf 0}$ (implied by (6)). The SETM is similar to the method used in Lancaster and Imbens, (1996) in spirit and can also be seen as an extension and refinement of the methods used in Qin, (1999) and Qin and Liang, (2011) as these methods are constrained to scalar $x$.

### v3

@@ Abstract
This study introduces a new approach to addressing the positive and unlabeled (PU) data through the double exponential tilting model (DETM) under a transfer learning framework. Traditional methods often fall short because they only apply to the common distributions (CD) PU data (also known as the selected completely at random PU data), where the labeled positive and unlabeled positive data are assumed to be from the same distribution. In contrast, our DETM’s dual structure effectively accommodates the more complex and underexplored different distribution (DD) PU data (also known as the selected at random PU data), where the labeled and unlabeled positive data can be from different distributions. We rigorously establish the theoretical foundations of DETM, including identifiability, parameter estimation, and asymptotic properties. Additionally, we move forward to statistical inference by developing a goodness-of-fit test for the CD assumption and constructing confidence intervals for the proportion of positive instances in the target domain. We leverage an approximated Bayes classifier for classification tasks, demonstrating DETM’s robust performance in prediction. Through theoretical insights and practical applications, this study highlights DETM as a comprehensive framework for addressing the challenges of PU data.
@@ 1.1 Positive and Unlabeled Data
Considering a binary response variable $Y\in\{0,1\}$ and features ${\bf X}\in\mathbb{R}^{p}$, positive and unlabeled (PU) data refer to datasets that consist of the following two samples: The first sample is labeled, but all observations in this sample have positive responses $y=1$; the second sample contains both positive ($y=1$) and negative ($y=0$) instances. However, the second sample is unlabeled, meaning we can only access the features ${\bf x}$ but have no access to their labels $y$. PU data are prevalent in many areas, including biomedical study, ecology, and machine learning.
@@
A case-control study is a research design widely used in biomedical research to investigate the causes of diseases. It compares two groups: a group of patients with the disease (the case group) and a group of healthy people (the control group). Data contamination is a common issue in case-control studies. For example, the underdiagnosis phenomenon refers to a situation where the false positive rate is very low, and the case group can generally be seen as pure and only contains patients. However, some medical conditions may go unrecognized in many patients, and they are misclassified as healthy ones. When underdiagnosis happens, many people in the control group are, in fact, patients. For example, Begg, (1994) and Godley and Schell, (1999) highlight an issue in a prostate cancer study, indicating that 20%-40% of the control group consists of latent prostate cancer cases, rendering the control group contaminated. In the underdiagnosis example, the case group can be seen as a positive-only sample, but the contaminated control group is an unlabeled sample containing both positive and negative cases because we do not know who are real healthy people and patients. In addition to biomedical studies, contaminated case-control data are also prevalent in econometrics (e.g., Lancaster and Imbens, 1996).
@@
Presence-only data can come from various fields, including ecology, epidemiology, and marketing. For example, ecological studies often rely on presence-only data to model a species’ habitat (Ward et al., 2009). The dataset comprises two samples: one sample contains confirmed presence locations of a species obtained through field surveys or other methods; the other sample contains some other locations sampled from the full study area. However, the presence status of the species in the latter sample is unknown. Thus, the first sample is positive only while the second sample contains both presence and non-presence locations but is unlabeled, meaning that we do not know which location has the confirmed presence of the species.
@@
PU data are commonly seen in many machine learning tasks. For instance, on an online video streaming platform, users click on videos that they find interesting. However, for videos that are not clicked, it is uncertain whether users are interested in them or not, as users may ignore those videos while browsing. In such a scenario, those clicked videos are positive-only data, while unclicked videos are considered unlabeled data for a recommendation system. In computer vision, tasks like facial recognition are often based on PU data. Take this specific example: We aim to identify a user from various images using some known reference pictures of that person. Here, the labeled sample comprises all clearly identified face images of that user. The unlabeled sample we wish to identify consists of photos of the user and other people. We refer interested readers to read Jaskie and Spanias, (2019) for more examples of PU data in machine learning.
@@ 1.2 Notation and Motivations
Even though much research has been done on PU data, most of it suffers from two limitations. The first limitation is that many existing works require the mixture proportion $\pi$ to be known. However, such an assumption is usually unrealistic as $\pi$ is rarely known.
The second and major limitation is that most existing research works are constrained to the common distributions (CD) PU data (also known as the selected at random PU data, see Bekker and Davis, 2020), which imposes the assumption that $p_{s}({\bf x}|y=1)=p_{t}({\bf x}|y=1)$. The CD assumption has been widely used for simplicity: one of the mixture components, $p_{t}({\bf x}|y=1)$, is readily identifiable because the positive-only data $\left\{({\bf x}_{i},y_{i}=1),i=1,\dots,n\right\}$ are from the same distribution as $p_{t}({\bf x}|y=1)$. Despite its popularity, the CD assumption does not always reflect the true underlying data-generating mechanism. In the contaminated case-control study example, it is reasonable to believe that the latent prostate cancer cases were misdiagnosed as healthy ones not only because of bad luck but also because their disease symptoms were not as severe (or typical) as patients in the case group (i.e., $p_{s}({\bf x}|y=1)\neq p_{t}({\bf x}|y=1)$).
As opposed to the CD assumption, the different distributions (DD) PU setting (also known as selected-at-random PU, see Bekker and Davis, 2020) allows $p_{s}({\bf x}|y=1)$ to be different from $p_{t}({\bf x}|y=1)$, thus it can better reflect the true data-generating mechanism. Figure 1 shows the difference between DD and CD PU data. However, despite its flexibility, the DD PU setting presents numerous challenges, the foremost being the identifiability of the distributions of unlabeled positive and unlabeled negative data (i.e., $p_{t}({\bf x}|y=1)$ and $p_{t}({\bf x}|y=0)$).
We aim to address the problem of statistical inference and classification on PU data by adopting the DD setting, which relaxes the often unrealistic CD assumption. A key advantage of our approach is that it does not require prior knowledge of the mixture proportion $\pi$.
Figure 1: We use $R=0$ (red) and $R=1$ (blue), respectively, to denote data from the target and source. In the CD setting, the labeled positive (blue triangles) and unlabeled positive (red triangles) are from the same distribution. However, they may be different in the DD setting.
@@ 1.3 Related Work and Challenges
PU data have been studied under different names in the statistical literature, including contaminated case-control data (e.g., Qin and Lawless, 1994; Lancaster and Imbens, 1996; Qin, 1999; Qin and Liang, 2011; Duan et al., 2020) and presence-only data (e.g., Ward et al., 2009; Song and Raskutti, 2019). However, the aforementioned works are constrained under the CD assumption and often rely on knowing the true value of the mixture proportion. One exception is Wang et al., (2021), where they consider the DD setting. However, they further require a smaller labeled verification sample from the target distribution. Thus, data from their setting cannot be seen as PU data.
In machine learning literature, most works on PU data focus on the CD assumption (e.g., Elkan and Noto, 2008; Ramaswamy et al., 2016; Garg et al., 2021). However, a few works try to address the DD setting. Kato et al., (2019) impose an “invariance of order” assumption, which states that, for observations in the labeled source data, the higher the probability $\Pr_{s}(y=1|{\bf x})$ is, the closer it is to the unlabeled positive distribution $p_{t}({\bf x}|y=1)$. In addition, their method requires the mixture proportion $\pi=\Pr_{t}(y=1)$ to be known. Bekker et al., (2019) propose a method similar to inverse probability weighting, which requires estimating the propensity score function. To ensure identifiability, they assume that the propensity score function needs to depend on fewer features than the predictive model (i.e., the shadow variable assumption in the missing data literature). Coudray et al., (2023) further establish risk bounds based on Bekker et al., (2019). Nevertheless, the shadow variables are generally unknown, let alone their existence, thus reducing the method’s applicability. Gong et al., (2021) follow a similar framework as in Bekker et al., (2019) but do not impose the shadow variable assumption. However, they require a highly restrictive assumption that the propensity score function has a value of less than 0.5 for all observations from the unlabeled data under the logistic regression model. Furmańczyk et al., (2022) use a similar idea as Bekker et al., (2019) and Gong et al., (2021), in which they prove the model identifiability under logistic regression. However, their prediction approach requires that the positive and negative instances in the unlabeled data are separable. In addition, existing works on the DD PU setting focus solely on classification (i.e., prediction) and do not address any statistical inference problems. Problems like the the goodness-of-fit of the CD condition, and inference of the mixture proportion $\pi$ are remained unexplored.
@@ 1.4 Our Contributions and Overview
Unlike the traditional propensity score framework used in existing literature (Bekker et al., 2019; Bekker and Davis, 2020), we treat the PU data based on a transfer learning framework where the data consist of two samples (a detailed comparison of the two frameworks is provided in Section 6). We further propose a new model, called the double exponential tilting model (DETM), for the DD PU data, and prove its identifiability. Based on the proposed model, we develop an Expectation-Maximization (EM) algorithm for estimating parameters using the empirical likelihood (EL) method. It is worth noting that our method works even if the mixture proportion $\pi$ is unknown. Lastly, we apply the proposed model to address inference and prediction tasks, including model selection, confidence intervals, hypothesis testing, and classification. The highlights of our contributions are summarized in the following.
  • Our proposed DETM, based on modeling the density ratios, is identifiable up to label switching. With some further mild assumptions, the model becomes fully identifiable.
@@
  • We investigate the asymptotic behavior of our proposed estimator and propose hypothesis testing procedures for the mixture proportion and the goodness-of-fit of the CD condition.
@@
The rest of this paper is organized as follows. Section 2 introduces the proposed DETM. Section 3 discusses how to estimate parameters under the proposed model. Section 4 establishes theoretical results for statistical inference and classification. Section 5 examines the empirical performance of the proposed methods using simulation studies and a real data application. We close the paper with concluding remarks given in Section 6. All the technical details and proofs are given in the supplementary materials.
@@ 2.1 Density Ratio is the Key
A commonly used strategy in distributional shift problems (or, broadly speaking, transfer learning) for connecting the source and target domains is to utilize the density ratio. For example, adjusting the covariate shift often relies on estimating the density ratio of the features ${\bf x}$ between the target and source domains (e.g., Shimodaira, 2000; Sugiyama et al., 2007), and adjusting the label shift often requires estimating the density ratio of the response between the target and source domains (e.g., Lipton et al., 2018; Tian et al., 2023; Kim et al., 2024).
The aforementioned works on transfer learning suggest that modeling the density ratios is a promising direction for solving the PU problem, which also involves a source and a target domain and belongs to the broad concept of transfer learning. In fact, examining the Bayes classifier for the target domain also reveals the key role of the density ratios. The Bayes classifier utilizes the posterior probability of the label and is optimal because it minimizes the prediction risk. The Bayes classifier on the target domain is defined as
@@
where $\mathbbm{1}(\cdot)$ is the indicator function. We can rewrite ${\Pr}_{t}(Y=1|{\bf X}={\bf x})$ using the Bayes rule as
@@
We can further write ${\Pr}_{t}(Y=1|{\bf X}={\bf x})$ by dividing both the numerator and denominator by $p_{s}({\bf x}|y=1)$:
@@
Equation (2) implies that the Bayes classifier hinges on the mixture proportion $\pi$ and two density ratios $\Lambda_{1}({\bf x})$ and $\Lambda_{2}({\bf x})$. This finding also points out that it is promising to model the DD PU data using the density ratios.
@@ 2.2 Identifiability Problem
We have underscored the importance of $\Lambda_{1}({\bf x})$ and $\Lambda_{2}({\bf x})$ in the previous section. In this section, we will show that if $\Lambda_{1}({\bf x})$ and $\Lambda_{2}({\bf x})$ are unconstrained, $\pi$, $\Lambda_{1}({\bf x})$, and $\Lambda_{2}({\bf x})$ are non-identifiable.
@@ 2.3 Double Exponential Tilting Model
Based on previous discussions, we use the ETM to connect the source and target domains. Rather than using an exponential tilting model for $p_{t}({\bf x},y)/p_{s}({\bf x},y)$ (e.g., Maity et al., 2023), we acknowledge that the target distribution $p_{t}({\bf x})$ is a mixture of $p_{t}({\bf x}|y=1)$ and $p_{t}({\bf x}|y=0)$. Thus, we connect the source sample distribution $p_{s}({\bf x}|y=1)$ with the two mixture components $p_{t}({\bf x}|y=1)$ and $p_{t}({\bf x}|y=0)$ using two ETMs separately. Thus, we name the proposed model as the DETM, where the details are given below.
@@
The remaining question is to address the identifiability of the DETM. As we have shown in (3), the key to model identifiability lies in $\pi\Lambda_{1}({\bf x})+(1-\pi)\Lambda_{2}({\bf x})$. In the next theorem, we prove that $\pi\Lambda_{1}({\bf x})+(1-\pi)\Lambda_{2}({\bf x})$ is “almost” identifiable and can become fully identifiable under some non-restrictive conditions. The proof is given in the supplementary materials.
@@
Under the DETM in (4), we model the density ratios as $\Lambda_{1}({\bf x})=\exp\left(\alpha_{1}+{\bf x}^{\rm T}{\boldsymbol{\beta}}_{{1}}\right)$ and $\Lambda_{2}({\bf x})=\exp(\alpha_{2}+{\bf x}^{\rm T}{\boldsymbol{\beta}}_{{2}})$, and we let $\pi={\Pr}_{t}(Y=1)$. If $(\pi,\alpha_{1},{\boldsymbol{\beta}}_{1},\alpha_{2},{\boldsymbol{\beta}}_{2})$ and $(\pi^{\ast},\alpha^{\ast}_{1},{\boldsymbol{\beta}}^{\ast}_{1},\alpha^{\ast}_{2},{\boldsymbol{\beta}}^{\ast}_{2})$ satisfy
@@
for any ${\bf x}\in\mathbb{R}^{p}$, then the two sets of parameters are either the same (i.e., $(\pi,\alpha_{1},{\boldsymbol{\beta}}_{1},\alpha_{2},{\boldsymbol{\beta}}_{2})=(\pi^{\ast},\alpha^{\ast}_{1},{\boldsymbol{\beta}}^{\ast}_{1},\alpha^{\ast}_{2},{\boldsymbol{\beta}}^{\ast}_{2})$) or subject to label switching: $(\pi,\alpha_{1},{\boldsymbol{\beta}}_{1},\alpha_{2},{\boldsymbol{\beta}}_{2})=(1-\pi^{\ast},\alpha^{\ast}_{2},{\boldsymbol{\beta}}^{\ast}_{2},\alpha^{\ast}_{1},{\boldsymbol{\beta}}^{\ast}_{1})$.
@@
Theorem 1 reveals that the DETM is identifiable up to exchanging the label $Y$. The label switching issue is inherent in the model specification in (4) due to the symmetry of $\Lambda_{1}({\bf x})$ and $\Lambda_{2}({\bf x})$. We can achieve full model identifiability by introducing some reasonable assumptions to solve the label-switching problem. For instance, if positive cases are less frequent in the unlabeled dataset (e.g., misdiagnosis is rare), we can impose the condition $\pi<0.5$, which immediately resolves the label-switching problem. In the rest of this paper, we assume the DETM is fully identifiable.
@@
The proposed DETM can be applied to both CD and DD PU data because CD data can be viewed as a special case of DD data with additional constraints. Under the CD assumption, the distribution of features of positive cases remains the same in the source and target domains. That is
@@
for any ${\bf x}$. We refer to the constrained model for CD PU data as the single exponential tilting model (SETM) because we only have the following model assumption:
@@
The SETM is a reduced model of DETM because we consider a reduced parameter space where $\alpha_{1}=0$ and ${\boldsymbol{\beta}}_{1}={\bf 0}$ (implied by (6)). The SETM is similar to the method used in Lancaster and Imbens, (1996) in spirit and can also be seen as an extension and refinement of the methods in Qin, (1999) and Qin and Liang, (2011) as these methods are constrained to scalar $x$.
Remark 3.
The DETM in (4) inherently assumes that $p_{s}(\mathbf{x}\mid y=1)$ dominates both $p_{t}(\mathbf{x}\mid y=1)$ and $p_{t}(\mathbf{x}\mid y=0)$, with all three distributions sharing a common support. In other words, this requires $\Lambda_{1}(\mathbf{x})>0$ and $\Lambda_{2}(\mathbf{x})>0$ on the common support. For instance, if $p_{t}(\mathbf{x}\mid y=0)$ and $p_{s}(\mathbf{x}\mid y=1)$ are perfectly separable, the assumption $\Lambda_{2}(\mathbf{x})>0$ will not hold, rendering our method inapplicable. In practice, the dominance assumption implies that data from $p_{t}(\mathbf{x}\mid y=0)$ (and similarly, data from $p_{t}(\mathbf{x}\mid y=1)$) should preferably overlap to some extent with data from $p_{s}(\mathbf{x}\mid y=1)$. In contrast, some methods rely on the separability assumption (e.g., Garg et al., 2022), require that the data are well-separated and may fail when significant overlap exists. It is important to emphasize that dealing with overlapping data is, in fact, more challenging. When $p_{t}(\mathbf{x}\mid y=1)$ is well-separated from other data, identifying examples from $p_{t}(\mathbf{x}\mid y=1)$ becomes straightforward. However, in overlapping scenarios, distinguishing between classes is considerably more difficult, highlighting the practical relevance of our approach.